OpenStudy (anonymous):

Pyramid A is a square pyramid with a base side length of 18 inches and a height of 9 inches. Pyramid B has a volume of 3136 cubic inches. How many times bigger is the volume of pyramid B than pyramid A? Give your answer as a percent.

OpenStudy (anonymous):

@paki @ParthKohli

OpenStudy (anonymous):

I got the volume of Pyramid A: Pyramid A's volume is 972 because V = Base side length^2 x h/3 V = 18^2 x 9/3 = 972.

OpenStudy (anonymous):

@sourwing @elementwielder @eliassaab

OpenStudy (anonymous):

divide the smaller number by the larger number then multiply that by 100 and round if needed

OpenStudy (anonymous):

972/3136 = 0.30994898 x 100 = 30.99. So would it be 31% or 30%?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

which one?

OpenStudy (anonymous):

go for the higher percentage since it is closer to the exact answer and sorry wasnt reading carefully enough

OpenStudy (anonymous):

31% is your final answer